Line/ goes through the point (5,2) and is perpendicular to the line through(-2,1) and (-4,-2). Find the equation (in slope form) for/.
Answer by mananth(16946)

slope of line passing through (-2,1) & (-4,-2)
x1 y1 x2 y2
-2 1 -4 -2

slope m = (y2-y1)/(x2-x1)
( -2 - 1 )/( -4 - -2 )
( -3 / -2 )
m= 1.50 OR 3/2
The line perpendicular to this line will have a slope of -2/3
Equation of line with slope -2/3 and passing through (5,2)
Find the intercept b
Plug value of the slope -2/3 and point (5,2) in
Y = m x + b
2.00 = -3.75 + b
b= 2 +3.75
b= 5.75
So the equation will be
Y = -0.75 x + 5.75
m.ananth@hotmail.ca